Consider the following balanced equation for the combustion of butane, a fuel often used in lighters.

2C

4

H

10

(g)+13O

2

(g)→8CO

2

(g)+10H

2

O(g)

2C4H10(g)+13O2(g)→8CO2(g)+10H2O(g) Complete the following table, showing the appropriate masses of reactants and products. If the mass of a reactant is provided, fill in the mass of other reactants required to completely react with the given mass, as well as the mass of each product formed. If the mass of a product is provided, fill in the required masses of each reactant to make that amount of product, as well as the mass of the other product that is formed.

Mass C

4

H

10

C4H10Mass O

2

O2Mass CO

2

CO2Mass H

2

O

H2O_____1.71 g

g__________4.72 g

g_________________________10.12 g

g____________________9.24 g

g222 mg

mg_________________________138 mg

mg_____


1
Expert's answer
2021-04-09T04:26:11-0400

The more butane is burnt, the more heat is produced. So calculating what quantites of butane undergo reaction in each case will show what amounts of heat are produced.

1) n(C4H10) = 1 mol

2) n(C4H10) = "{\frac {2} {13}}" mol accroding to equation, as n(O2) = 1 mol and n(C4H10) = "{\frac {2} {13}}"n(O2)

3) n(C4H10) = "{\frac {1} {4}}" mol according to equation, as n(CO2 produced) = 1 mol and n(C4H10) = "{\frac {2} {8}}"n(CO2)

4) n(C4H10) = "{\frac {1} {5}}" mol according to equation, as n(H2O produced) = 1 mol and n(C4H10) = "{\frac {2} {10}}"n(H2O)

the least amount of butane involved is in the second case, so the least amount of heat generated in